# When Can Transformers Count to n?

### v1

@@ Abstract
Large language models based on the transformer architectures can solve highly complex tasks. But are there simple tasks that such models cannot solve? Here we focus on very simple counting tasks, that involve counting how many times a token in the vocabulary have appeared in a string. We show that if the dimension of the transformer state is linear in the context length, this task can be solved. However, the solution we propose does not scale beyond this limit, and we provide theoretical arguments for why it is likely impossible for a size limited transformer to implement this task. Our empirical results demonstrate the same phase-transition in performance, as anticipated by the theoretical argument. Our results demonstrate the importance of understanding how transformers can solve simple tasks.
@@ 1 Introduction
Large language models (LLMs) have demonstrated striking performance on a wide array of tasks, from creative writing to solving complex math problems. Given these successes a key question then arises: what can these models do, and just as importantly what can they not do. There are multiple ways to address this question of expressiveness of LLMs. First, it can be studied empirically by probing LLMs and seeking relatively simple tasks that they cannot perform successfully. Indeed recent work has found several such tasks, including “needle in haystack” [7, 6] as well as extrapolating tasks to longer sequences [8]. A second, complementary, approach is theoretical studies which chart the computational capabilities of transformers [12, 17].
@@
Our interest in this problem is that it seems like a basic module needed for working with long input streams. Indeed, there is a long line of work on sketching and streaming, that studies similar tasks [1, 4]. Furthermore, this task can be viewed as an extension of the “needle in haystack” problem which has recently garnered much attention, since it was shown that most LLMs struggle with it for long context. Specifically, in the needle in haystack problem, the goal is to find a specific string in a long text. In the count problem, the model is tasked with counting how many times a given string has appeared, which is a harder task than finding one appearance.
@@
There is, however, a key difference between the needle in haystack problem and the counting problem we consider above. The needle in haystack problem is clearly solvable by transformers, regardless of the context length. This is because detecting a similar token and extracting it (or nearby tokens) is a simple task for single attention head [e.g., using induction heads 10]. On the other hand, for the counting problem, as we argue here, it is much less clear that transformers can solve it for arbitrary context length. We remark that we write that transformers cannot perform a task, we are referring to transformers whose number of parameters is not dependent on the context size.
@@
Modern LLMs indeed struggle with counting tasks (see Sec. 7). Of course if these models can use code, the task becomes easy, but our focus is on understanding the capabilities of the transformer architecture itself. Specifically, we ask whether transformers have an architectural limitation related to the counting task.
Considering why transformers could struggle with this task, a key factor is the fact the averaging nature of the softmax attention. Intuitively, a simple way to solve the counting task is to let the query token attend to all previous token and assign high attention weights to those that are identical to it, and low to the others. This is indeed possible via the Q/K/V matrices. However, the attention mechanism then normalizes these weights, so that their sum is one, regardless of the count of the query token in the sequence. As we argue in Sec. 4, this already provides some insight into the limitations of transformers for this task.
@@
We next turn to ask when transformers can count. We show a construction that works as long as the transformer embedding size $d$ is greater than the vocabulary size $m$. Our construction uses a one-hot embedding, or more generally an orthogonal embedding, of the vocabulary, which allows the model to maintain a histogram of counts of tokens previously observed. When $d<m$ this orthogonal construction is no longer possible. A natural approach is to consider the “most orthogonal possible” embedding (a notion formalized in Welch bounds [19]), and try to use it within a similar scheme. However, we show that this does not allow implementing the histogram solution beyond $d>m$.
@@
The above discussion suggests that in the $d<m$ regime, the naive approach of a histogram does not seem to work. Our study of this regime reveals both a positive and a negative result. On the positive side, we show that there does exist a construction that allows counting, which can be done with a single transformer layer. On the negative side, we prove that this construction requires an MLP width that grows with context size, meaning it is not applicable to arbitrarily long contexts. Indeed, when optimizing transformers we find that they fail to learn in this regime of $d>m$.
@@
We next study a somewhat more complex counting task, which we refer to as “Most Frequent Element”. Here we present the model with a sequence of tokens, and ask for the count of the most frequent token. This is the same as taking the maximum of the histogram of the counts. Similarly to Query Count, we show that in this case, a solution exists for $d<m$ based on an orthogonal construction. However, for $d>m$ we show, using a communication complexity argument, that no solution exists for a one layer transformer. Thus again we obtain a phase transition for counting at $d=m$.
@@
Our results highlight the importance of studying basic counting problems, and its dependence on vocabulary size. They also point to limits of solving seemingly simple problems using transformers, and further emphasize the advantages of using code as a tool to sidestep these issues.
@@ 2 Related Work
Since the introduction of transformer architectures [16] and the success of LLMs, there has been much work on evaluating their performance on various tasks [e.g., see 14]. In particular, much recent work has explored performance on long context tasks, the dependence of accuracy on context length [8], and the ability of models to extrapolate to lengths beyond those seen at training [2].
The fact that models often do not perform well on these evaluations has prompted works that try to pinpoint the inherent limitations of transformer models. One line of work is to use computational complexity lower bound approaches to show that transformers of a certain size and architecture cannot implement a given function of interest. For example, Sanford et al. [12] show that a certain functions cannot be implemented without transformer size scaling with input size.
A related line of work is to relate transformers to known complexity classes. For example it has been shown that transformers are Turing complete [17], and that transformers with bounded percision can only solve problems in uniform $TC^{0}$ [9]. Chain-of-Thought [18] has also been analyzed from an expressiveness viewpoint, demonstrating it can substantially improve the expressive power of transformers [5].
@@
(b) Solving QC using CountAttend: This solution works for all $d$, but requires an MLP for inverting numbers, and this MLP is of size $n$ (which can be prohibitive). To count the number of tokens with $x_{i}=4$, the last token attends to the others such that only tokens with $x_{i}=4$ receive large weights. This results in weights that are non-zero only for $x_{i}=4$, and the resulting weight on these is the inverse of the count of $4$ (ie 0.5 in this case). Then this inverse is moved to the last element of the value vector, using a positional embedding coordinate that is $1$ only for last token $n$. Finally, the inverse count needs to be inverted to get the desired count, and this requires the “Inversion MLP”.
@@ 3 Problem Setup
  • The query, key, value matrices for layer $i$, attention head $j$ will be $Q_{i,j},K_{i,j},V_{i,j}$. All are matrices in $\mathbb{R}^{dh\times d}$.
@@
Most of our solutions work with an architecture consisting of a single layer and a single head. When this is the case we omit the indices $i$ and/or $j$, respectively from our notation. Also note that if $h=1$ then $D=d$.
4 The Need for Positional Embeddings
Before analyzing specific solutions for counting, we remark on a general limitation of transformers for the counting problems we consider. [^1] We note that a similar observation was concurrently made in [3].
Transformers use self attention to average over previous representations. The fact that they average rather than sum leads to an interesting limitation on their ability to count. Specifically, it is easy to see that for variable context size, they cannot perform any counting task without the use of positional embedding. Consider the QC task and an input sequence $S_{1}=x_{1},\ldots,x_{n}$, where the goal is to return the count of $x_{n}$ in the sequence. Now consider the length $2n$ sequence $S_{2}=x_{1},\ldots,x_{n},x_{1},\ldots,x_{n}$. The correct output for this sequence is twice the correct output for $S_{1}$. However, a transformer without positional embeddings that is applied to $S_{1}$ will have exactly the same output as the one for $S_{2}$. This follows because averaging is invariant to input duplication.
The above restriction no longer holds when positional embeddings are used, and it is easy to see that it can be rectified with even a simple positional embedding that just signifies the last position (see our construction in Sec. 5.3). This implies that if a transformer has access to the legnth of the sequence, it may make it easier to count.
Another thing to note is that while the above difficulty arises for counting, it does not arise if we are interested in calculating proportions (e.g., what is the fraction of the items of the sequence that are equal to $x_{n}$).
@@ 5 Analyzing Query Count (QC)
In this section we focus on the QC problem, and ask which transformer architectures can implement it successfully. We first show in Sec. 5.1 that if $d>2m$ a one-head one-layer transformer can implement QC. We refer to this as the histogram solution. We then show that the histogram solution stops working if $d<m$. The natural question is then whether there are other solutions for the $d<m$ case. We argue that in this case solutions are likely to require calculating the function $1/x$, and show that this function would require an MLP layer of width $n^{2}$. This means we cannot expect the transformer to extrapolate to long context sizes, and therefore a one-layer transformer is unlikely to be able to implement QC.
@@ 5.1 A “Histogram” Solution for $d>2m$
For the Query Count problem, if $d>2m$, there exists a transformer that solves it, which has one layer, one head, and an MLP with $d$ neurons.
@@
where $v_{i}$ is the embedding into $\mathbb{R}^{d}$ of the dictionary item $i$. This is possible because of the assumption $d>m$. For simplicity, we assume that $\boldsymbol{v}_{i}=\boldsymbol{e}_{i}$ where $\boldsymbol{e}_{i}$ is the standard basis in $\mathbb{R}^{d}$.
@@
Remarks: 1) Note that the above solution will only work for input of a fixed length $n$. As noted in 4, a transformer without positional embeddings (as the one in our construction) cannot possibly count inputs of variable lengths. Our construction here can be extended to variable lengths by using positional embeddings, together with an MLP for computing $1/x$. We elaborate on such an approach in the next section.
2) We can also implement the above histogram scheme with one head, by taking advantage of the residual connections. The idea is to use half of the coordinates in the embedding dimension to store the result of the attention module, and pass the original token to the MLP using the residual connections.
3) In case $v_{1},\ldots,v_{m}$ are orthonormal but its not necessarily the case that $v_{i}=e_{i}$ (i.e. they are not one-hot), the MLP will have to take the dot product of $n*u_{1}$ and $u_{2}$ to extract the count. This is less natural to do with RELU gates. One could, however, first apply to $u_{1}$ and $u_{2}$ a linear transformation (a rotation basically), that changes the basis to the standard basis and then extract the count as before.
@@ 5.2 The Histogram Solution Breaks for $d<m$
Our solution in the previous section uses the fact that if $d>m$ we can embed the dictionary into orthogonal vectors in $\mathbb{R}^{d}$. When $d<m$ this is not possible. One may try to extend this solution by embedding the dictionary into a collection of “almost” orthogonal vectors. However any collection of $m$, say $m\geq 2d$, vectors in $\mathbb{R}^{d}$, contains a pair of vectors whose inner product is at least $\Omega(1/\sqrt{d})$ in absolute value. This is a result of the Welch bounds [19] which provide upper bounds on the minimum dot product between $m$ unit-vectors in $\mathbb{R}^{d}$. This implies the following lower bound, which states that counting will fail in this regime.
@@
Consider the “Histogram” solution for the counting problem presented in Sec. 5.1, and embedding vectors $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}$. For an input $\bar{\mathbf{x}}=(x_{1},\dots,x_{n})$ to the counting problem, denote by $c_{x_{n}}$ the correct solution and by $\texttt{hist}(\bar{\mathbf{x}})$ the output of the “Histogram” solution. If $m\geq 2d$, then for any embedding vectors $v_{i}$’s there are inputs to the counting problem for which: $|\texttt{hist}(\bar{\mathbf{x}})-c_{x_{n}}|\geq 0.25\sqrt{n}$.
@@
Let $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}\in\mathbb{R}^{d}$ with $m\geq 2d$, and let $A=\max_{i\neq j}|\langle\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{j}\rangle|$. Assume without loss of generality that $A=\boldsymbol{v}_{1}\cdot\boldsymbol{v}_{2}$. By the Welch bounds [19] for $k=1$ we have that $A\geq\frac{1}{\sqrt{2d-1}}$. Consider the input $x_{1},\dots,x_{n}$ to the counting problem where $x_{1},\dots,x_{n-c}$ are equal to the same token which is different from $x_{n}$ and mapped to the embedding $v_{1}$, and $x_{n-c},\dots,x_{n}$ are all equal to $x_{n}$ which is mapped to embedding $\boldsymbol{v}_{2}$. Then the output for the histogram solution is:
@@
The theorem implies that even if the dictionary size is only linear in the embedding dimension, any solution will incur an error that grows with the context size. In practice, the dictionary can contain millions of tokens, while the embedding dimension is at most a few thousands, thus this error can be quite large. Note that picking the embedding vectors at random (e.g. i.i.d Gaussians) will only suffer an even greater error than what is stated in the theorem, since the inner product between each two vectors will be larger (with high probability) than the lower bound we used in the proof.

### v2

@@ Abstract
Large language models based on the transformer architecture can solve highly complex tasks. But are there simple tasks that such models cannot solve? Here we focus on very simple counting tasks, that involve counting how many times tokens in the vocabulary appeared in a string. We show that if the dimension of the transformer state is linear in the context length, this task can be solved. However, the solution we propose does not scale beyond this limit, and we provide theoretical arguments for why it is likely impossible for a size-limited transformer to implement this task. Our empirical results demonstrate the same phase-transition in performance, as anticipated by the theoretical argument. Our results demonstrate the importance of understanding how transformers can solve simple tasks.
@@ 1 Introduction
Large language models (LLMs) have demonstrated striking performance on a wide array of tasks, from creative writing to solving complex math problems. Given these successes, a key question arises: what can these models do, and just as importantly what can they not do. There are multiple ways to address this question of expressiveness of LLMs. First, it can be studied empirically by probing LLMs and seeking relatively simple tasks that they cannot perform successfully. Indeed recent work has found several such tasks, including “needle in haystack” [7, 6] as well as extrapolating tasks to longer sequences [8]. A second, complementary, approach is theoretical studies which chart the computational capabilities of transformers [13, 18].
@@
Our interest in this problem is that it seems like a basic module needed for working with long input streams. Indeed, there is a long line of work on sketching and streaming, that studies similar tasks [1, 4]. Furthermore, this task can be viewed as an extension of the “needle in haystack” problem which has recently garnered much attention, since it was shown that most LLMs struggle with it for long context. In the needle in haystack problem, the goal is to find a specific string in a long text. In our counting problem, the model is tasked with counting how many times a given string has appeared, which is a harder task than finding one appearance.
@@
There is, however, a key difference between the needle in haystack problem and the counting problem we consider above. The needle in haystack problem is clearly solvable by transformers, regardless of the context length. This is because detecting a similar token and extracting it (or nearby tokens) is a simple task for single attention head [e.g., using induction heads 10]. On the other hand, for the counting problem, as we argue here, it is much less clear that transformers can solve it for arbitrary context length. Note that when we write that transformers cannot perform a task, we are referring to transformers whose number of parameters is not dependent on context size.
@@
Modern LLMs indeed struggle with counting tasks (e.g., see Sec. 6 and [3]). Of course if these models can use code, the task becomes easy, but our focus is on understanding the capabilities of the transformer architecture itself. Specifically, we ask whether transformers have an architectural limitation related to the counting task.
@@
We next turn to ask when transformers can count. We show a construction that works as long as the transformer embedding size $d$ is greater than the vocabulary size $m$. Our construction uses one-hot embeddings, or more generally orthogonal embeddings, of the vocabulary, which allow the model to maintain a histogram of counts of tokens previously observed. When $d<m$ this orthogonal construction is no longer possible. A natural approach is to consider the “most orthogonal possible” embedding (a notion formalized in Welch bounds [20]), and try to use it in a similar scheme. However, we show that this does not allow implementing the histogram solution if the dimension is smaller than the vocabulary.
@@
The above discussion suggests that in the $d<m$ regime, [^1] Some of our results involve regimes such $d<cm$ for a small constant $c$. We leave $c$ out for simplicity. the naive histogram approach does not seem to work. Our study of this regime reveals both a positive and a negative result. On the positive side, we show that there does exist a construction that allows counting, which can be done with a single transformer layer. On the negative side, we prove that this construction requires an MLP width that grows with context size, meaning it is not applicable to arbitrarily long contexts. Indeed, when training transformers on these counting tasks, we find that they fail in the $d<m$ regime.
@@
We next study a somewhat more complex counting task, which we refer to as “Most Frequent Element”. Here we present the model with a sequence of tokens, and ask for the count of the most frequent token. This is the same as taking the maximum of the histogram of the counts. Similarly to the Query Counting task, we show that in this case, a solution exists for $d>m$ based on an orthogonal construction. However, for $d<m$ we show, using a communication complexity argument, that no solution exists for a one layer transformer. Thus again we obtain a phase transition for counting at $d=m$.
@@
Our results highlight the importance of studying basic counting problems, and their dependence on vocabulary size. They also point to limits of solving seemingly simple problems using transformers, and further emphasize the advantages of using code as a tool to sidestep these issues.
@@ 2 Related Work
Since the introduction of transformer architectures [17] and the success of LLMs, there has been much work on evaluating their performance on various tasks [e.g., see 15]. In particular, much recent work has explored performance on long context tasks, the dependence of accuracy on context length [8], and the ability of models to extrapolate to lengths beyond those seen at training [2].
The fact that models often do not perform well on these evaluations has prompted works that try to pinpoint the inherent limitations of transformer models. One line of work is to use computational complexity lower bound approaches to show that transformers of a certain size and architecture cannot implement a given function of interest. For example, Sanford et al. [13] show that certain functions cannot be implemented without the transformer size scaling with input size. A similar limitation is also shown in [11] for certain compositions of functions.
A related line of work is to relate transformers to known complexity classes. For example it has been shown that transformers can efficiently approximate Turing machines [18], and that transformers with bounded percision can only solve problems in uniform $TC^{0}$ [9]. Chain-of-Thought [19] has also been analyzed from an expressiveness viewpoint, demonstrating it can substantially improve the expressive power of transformers [5].
@@
(b) Solving QC using CountAttend: this solution works for all $d$, but requires an MLP for inverting numbers, and we show that this MLP need to be of size $n$ (which can be prohibitive). To count the number of tokens with $x_{i}=4$, the last token attends to the others such that only tokens with $x_{i}=4$ receive large weights. This results in weights that are non-zero only for $x_{i}=4$, and the resulting weight on these is the inverse of the count of $4$ (i.e., 0.5 in this case). Then this inverse is moved to the last element of the value vector, using a positional embedding coordinate that is $1$ only for last token $n$. Finally, the inverse count needs to be inverted to get the desired count, and this requires the “Inversion MLP”.
@@ 3 Problem Setup
  • The query, key, value matrices for layer $i$, attention head $j$ are denoted by $Q_{i,j},K_{i,j},V_{i,j}$. All are matrices in $\mathbb{R}^{dh\times d}$.
@@
Most of our solutions work with an architecture consisting of a single layer and a single head. When this is the case we omit the indices $i$ and $j$ from our notation. Also note that if $h=1$ then $D=d$. In our theoretical results we do not use normalization layers, although we can easily add degenerate normalization that doesn’t alter the input, as was done in [13]. In our empirical results we use standard transformer layers that do include normalization layers.
@@ 4 Analyzing Query Count (QC)
In this section we focus on the QC problem, and ask which transformer architectures can implement it successfully. We first remark on a general limitation of transformers without positional embeddings for the counting problems we consider. [^2] We note that a similar observation was made in [3], we provide it here for completeness. We next show in Sec. 4.2 that if $d>2m$ a one-head one-layer transformer can implement QC. We refer to this as the histogram solution. We then show that the histogram solution stops working if $d<m$. The natural question is then whether there are other solutions for the $d<m$ case. We argue that in this case, solutions are likely to require calculating the function $1/x$, and show that this function would require an MLP layer of width $n$. This means we cannot expect the transformer to extrapolate to long context sizes, and therefore a one-layer transformer is unlikely to be able to implement QC, at least using the two most natural solutions.
4.1 The Need for Positional Embeddings
Transformers use self attention to average over previous representations. The fact that they average, rather than sum, leads to an interesting limitation on their ability to count. Specifically, it is easy to see that for variable context size, they cannot perform any counting task without the use of positional embedding. Consider the QC task and an input sequence $S_{1}=x_{1},\ldots,x_{n}$, where the goal is to return the count of $x_{n}$ in the sequence. Now consider the length $2n$ sequence $S_{2}=x_{1},\ldots,x_{n},x_{1},\ldots,x_{n}$. The correct output for this sequence is twice the correct output for $S_{1}$. However, a transformer without positional embeddings that is applied to $S_{1}$ will have exactly the same output as the one for $S_{2}$. This follows because averaging is invariant to input duplication.
The above restriction no longer holds when positional embeddings are used, and it is easy to see that it can be rectified with even a simple positional embedding that just signifies the last position (see our construction in Sec. 4.4). This implies that if a transformer has access to the legnth of the sequence, it may make it easier to count. Another thing to note is that while the above difficulty arises for counting, it does not arise if we are interested in calculating proportions (e.g., what is the fraction of the items of the sequence that are equal to $x_{n}$).
@@ 4.2 A “Histogram” Solution for $d>2m$
For the Query Count problem and any context length $n>0$, if $d>2m$, there exists a transformer that solves it, which has one layer, one head, and an MLP with $d$ neurons.
@@
where $\boldsymbol{v}_{i}$ is the embedding into $\mathbb{R}^{d}$ of the dictionary item $i$. This is possible because of the assumption $d>m$. For simplicity, we assume that $\boldsymbol{v}_{i}=\boldsymbol{e}_{i}$ where $\boldsymbol{e}_{i}$ is the standard basis in $\mathbb{R}^{d}$.
@@
Remarks: 1) Note that the above solution will only work for input of a fixed length $n$. As noted in 4.1, a transformer without positional embeddings (as the one in our construction) cannot possibly count inputs of variable lengths. Our construction here can be extended to variable lengths by using positional embeddings, together with an MLP for computing $1/x$. We elaborate on such an approach in the next section.
2) We can also implement the above histogram scheme with one head, by taking advantage of a residual connection. The idea is to use half of the coordinates in the embedding dimension to store the result of the attention module, and pass the original token to the MLP using the residual connection.
3) In the construction above, we assumed that $\boldsymbol{v}_{i}$ are the standard basis. However, a similar construction is possible when $\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{m}$ are orthonormal, but not one-hot. In this case, the MLP will have to take the dot product of $n*\boldsymbol{u}_{1}$ and $\boldsymbol{u}_{2}$ to extract the count. This is less natural to do with RELU gates. One could, however, first apply to $\boldsymbol{u}_{1}$ and $\boldsymbol{u}_{2}$ a linear transformation (i.e., a rotation), that changes the basis to the standard basis and then extract the count as before.
@@ 4.3 The Histogram Solution Breaks for $d<m$
Our solution in the previous section uses the fact that if $d>m$ we can embed the dictionary into orthogonal vectors in $\mathbb{R}^{d}$. When $d<m$ this is not possible. One may try to extend this solution by embedding the dictionary into a collection of “almost” orthogonal vectors. However any collection of $m$ $m\geq 2d$ vectors in $\mathbb{R}^{d}$ contains a pair of vectors whose inner product is at least $\Omega(1/\sqrt{d})$ in absolute value. This is a result of the Welch bounds [20] which provide upper bounds on the minimum dot product between $m$ unit-vectors in $\mathbb{R}^{d}$. This implies the following lower bound, which states that counting will fail in this regime.
@@
Consider the “Histogram” solution for the counting problem presented in Sec. 4.2, and embedding vectors $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}$. For an input $\bar{\mathbf{x}}=(x_{1},\dots,x_{n})$ to the counting problem, denote by $c_{x_{n}}$ the correct solution and by $\texttt{hist}(\bar{\mathbf{x}})$ the output of the “Histogram” solution. [^3] The Histogram solution in this case is $\boldsymbol{v}_{x_{n}}\cdot\sum_{j}c_{j}\boldsymbol{v}_{j}$, which is the natural generalization to the non-orthogonal case. As noted in Remark 3, it is generally more elaborate to implement because of the dot product, but as Theorem 4.2 shows, this solution has an inherent limitation, irrespective of this implementation difficulty. If $m\geq 2d$, then for any embedding vectors $\boldsymbol{v}_{i}$’s there are inputs to the counting problem for which: $|\texttt{hist}(\bar{\mathbf{x}})-c_{x_{n}}|\geq 0.25\sqrt{n}$.
@@
Let $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}\in\mathbb{R}^{d}$ with $m\geq 2d$, and let $A=\max_{i\neq j}|\langle\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{j}\rangle|$. Assume without loss of generality that $A=\boldsymbol{v}_{1}\cdot\boldsymbol{v}_{2}$. By the Welch bounds [20] for $k=1$ we have that $A\geq\frac{1}{\sqrt{2d-1}}$. Consider the input $x_{1},\dots,x_{n}$ to the counting problem where $x_{1},\dots,x_{n-c}$ are equal to the same token which is different from $x_{n}$ and mapped to the embedding $\boldsymbol{v}_{1}$, and $x_{n-c},\dots,x_{n}$ are all equal to $x_{n}$ which is mapped to embedding $\boldsymbol{v}_{2}$. Then the output for the histogram solution is:
@@
The theorem implies that even if the dictionary size is only linear in the embedding dimension, any solution will incur an error that grows with the context size. In practice, the dictionary can contain millions of tokens, while the embedding dimension is at most a few thousands, thus this error can be quite large. Note that picking the embedding vectors at random (e.g. i.i.d Gaussians) will result in an even greater error than what is stated in the theorem, since the inner product between each two vectors will be larger (with high probability) than the lower bound we used in the proof.